# The Uniform Mordell–Lang Conjecture

### v3

@@ 4.1. A result in family
Let $\mathbf{H}^{\circ}_{\mathrm{gen}}$ be the subset of $\mathbf{H}^{\circ}(\overline{\mathbb{Q}})$ as defined by (3.9). If the stabilizer of $X$ in $A$ (denoted by $\mathrm{Stab}_{A}(X)$) has positive dimension, then $X^{\circ}=\emptyset$ by definition of Ueno locus and hence the proposition trivially holds true. Thus we may and do assume that $\mathrm{Stab}_{A}(X)$ is finite. Then the point $s\in\mathbf{H}^{\circ}(\overline{\mathbb{Q}})$ which parametrizes $(X,(A,L))$ is in $\mathbf{H}_{\mathrm{gen}}^{\circ}$. As above, we identify $X=\mathscr{X}_{s}$ and $A=\mathcal{A}_{s}$.
Apply Proposition 4.2 to $S=\mathbf{H}^{\circ}$. Then we obtain constants $c_{1}^{\prime}$, $c_{2}^{\prime}$ and $c_{3}^{\prime}$ depending only on $r$, $8^{r}d$ and $\mathbf{H}^{\circ}$ (and hence only on $g$, $d$ and $r$). The set $\Sigma_{s}:=\left\{x\in\mathscr{X}_{s}^{\circ}(\overline{\mathbb{Q}})=X^{\circ}(\overline{\mathbb{Q}}):\hat{h}_{{\mathcal{L}}|_{A}}(x)\leq c_{1}^{\prime}\max\{1,h_{\overline{\mathbf{H}^{\circ}},{\mathcal{M}}}(s)\}-c_{3}^{\prime}\right\}$ is contains in $X^{\prime}$, for some proper Zariski closed $X^{\prime}\subsetneq X$ with $\deg_{{\mathcal{L}}|_{A}}(X^{\prime})<c_{2}^{\prime}$.
Notice that $\deg_{L}(X^{\prime})\leq\deg_{{\mathcal{L}}|_{A}}(X^{\prime})$ since $8c_{1}(L)=c_{1}({\mathcal{L}}|_{A})$, and hence $\deg_{L}(X^{\prime})<c_{2}^{\prime}$.
@@
Hence we are done with $c_{1}^{\prime}$ replaced by $c_{1}^{\prime}/4(c(g)+c^{\prime}(g))$ and $c_{3}^{\prime}$ replaced by $(c_{3}^{\prime}+2c_{1}^{\prime})/4$. ∎
@@
Let $(A,L)$ be a polarized abelian variety with $\deg_{L}A=l$ and let $X$ be an irreducible subvariety with $\deg_{L}X=d$ such that $X$ is not contained in any proper subgroup of $A$.
By Lemma 2.2.(iv), there exist a principally polarized abelian variety $(A_{0},L_{0})$ defined over $\overline{\mathbb{Q}}$ and an isogeny $u_{0}\colon A\rightarrow A_{0}$ such that $L=u_{0}^{*}L_{0}$; moreover, $\deg(u_{0})=l/g!$. A basic property of Faltings heights [Ray85, Prop.1.4.1] is
@@
It is well known that there exists an isogeny $u\colon A_{0}\rightarrow A$ such that $u_{0}\circ u=[\deg u_{0}]$ on $A_{0}$. So $u^{*}L=(u_{0}\circ u)^{*}L_{0}=[l/g!]^{*}L_{0}$, and $\deg u=(\deg u_{0})^{2g-1}=(l/g!)^{2g-1}$.
Let $X_{0}$ be an irreducible component of $u^{-1}(X)$. Then $u(X_{0})=X$. Moreover, $X_{0}$ is not contained in any proper subgroup of $A_{0}$.
By the projection formula for intersection numbers, we have
(4.14) $d^{\prime}:=\deg_{u^{*}L}X_{0}=c_{1}(u^{*}L)^{g-\dim X}\cdot[X_{0}]=\deg(u|_{X_{0}})c_{1}(L)^{g-\dim X}\cdot[X]\leq\deg(u)\deg_{L}X=d(l/g!)^{2g-1}.$
By the definition of Ueno locus, we have $X^{\circ}=u(X_{0})^{\circ}\subseteq u(X_{0}^{\circ})$.
Let $c_{0,1}^{\prime}(g,r,d^{\prime})>0$, $c_{0,2}^{\prime}(g,r,d^{\prime})>0$ and $c_{0,3}^{\prime}(g,r,d^{\prime})>0$ be the constants from Proposition 4.4 with $d$ replaced by $d^{\prime}$. Set $c_{1}^{\prime}:=\min_{1\leq d^{\prime}\leq d(l/g!)^{2g-1}}\{c_{0,1}^{\prime}(g,r,d^{\prime})\}>0$, $c_{2}^{\prime}:=\max_{1\leq d^{\prime}\leq d(l/g!)^{2g-1}}\{c_{0,2}^{\prime}(g,r,d^{\prime})\}>0$ and $c_{3}^{\prime}:=\max_{1\leq d^{\prime}\leq d(l/g!)^{2g-1}}\{c_{0,3}^{\prime}(g,r,d^{\prime})\}>0$. Then $c_{1}^{\prime}$, $c_{2}^{\prime}$ and $c_{3}^{\prime}$ depend only on $g$, $r$ and $d$. Proposition 4.4 yields the following assertion: the set $\Sigma_{0}:=\left\{x_{0}\in X_{0}^{\circ}(\overline{\mathbb{Q}}):\hat{h}_{L_{0}\otimes(L_{0})_{-}}(x_{0})\leq c_{1}^{\prime}\max\{1,h_{\mathrm{Fal}}(A_{0})\}-c_{3}^{\prime}\right\}$ is contained in $X_{0}^{\prime}(\overline{\mathbb{Q}})$ for some Zariski closed irreducible $X_{0}^{\prime}\subseteq X_{0}$ with $\deg_{L_{0}}(X^{\prime})\leq c_{2}^{\prime}$.
Let $\Sigma=\left\{x\in X^{\circ}(\overline{\mathbb{Q}}):(g!/l)^{2}\hat{h}_{L\otimes L_{-}}(x)\leq c_{1}^{\prime}\max\{1,h_{\mathrm{Fal}}(A)\}-c_{3}^{\prime}-(1/2)\log(l/g!)\right\}$. As $X^{\circ}\subseteq u(X_{0}^{\circ})$ and $u^{*}(L\otimes L_{-})\cong(L_{0}\otimes(L_{0})_{-})^{\otimes(l/g!)^{2}}$ (recall that $u^{*}L=[l/g!]^{*}L_{0}$), (4.13) yields $\Sigma\subseteq u(\Sigma_{0})$.
Let $X^{\prime}:=u(X_{0}^{\prime})$. Then $X^{\prime}$ is proper Zariski closed in $X$, $\Sigma\subseteq X^{\prime}(\overline{\mathbb{Q}})$, and
$\deg_{L}X^{\prime}$ $=c_{1}(L)^{g-\dim X^{\prime}}\cdot[X^{\prime}]$
$\leq c_{1}(u^{*}L)^{g-\dim X^{\prime}}\cdot[X_{0}^{\prime}]$
$=\deg_{u^{*}L}X_{0}^{\prime}\leq c_{2}^{\prime}.$
Replace $c_{1}^{\prime}$ by $(l/g!)^{2}c_{1}^{\prime}$ and $c_{3}^{\prime}$ by $(l/g!)^{2}(c_{3}^{\prime}+(1/2)\log(l/g!))$. Then Proposition 4.1 holds true.
@@ 5. Applying equidistribution
Let $g$, $l$, $r$ and $d$ be positive integers. There exist constants $c_{2}^{\prime\prime}=c_{2}^{\prime\prime}(g,l,r,d)>0$ and $c_{3}^{\prime\prime}=c_{3}^{\prime\prime}(g,l,r,d)>0$ satisfying the following property. For
  • and each irreducible subvariety $X$ of $A$, defined over $\overline{\mathbb{Q}}$, with $\dim X=r$ and $\deg_{L}X=d$, such that $X$ generates $A$,
@@
(5.1) $\Sigma:=\left\{x\in X^{\circ}(\overline{\mathbb{Q}}):\hat{h}_{L\otimes L_{-}}(x)\leq c_{3}^{\prime\prime}\right\},$
where $L_{-}=[-1]^{*}L$, is contained in $X^{\prime}(\overline{\mathbb{Q}})$, for some proper Zariski closed $X^{\prime}\subsetneq X$ with $\deg_{L}(X^{\prime})<c_{2}^{\prime\prime}$.
The basic idea to prove this proposition is similar to the proof of Proposition 4.1, i.e. to put all $(A,L)$ and $X$ into (finitely many) families over the components of the restricted Hilbert scheme from (3.6). The core of this section is the result in family Proposition 5.2.
5.1. A result in family
@@
For each $1\leq r\leq g$ and each $d\geq 1$, let $\mathbf{H}^{\circ}:=\mathbf{H}^{\circ}_{r,d}(\mathfrak{A}_{g}/\mathbb{A}_{g})$ be the restricted Hilbert scheme from (3.6); it is a variety defined over $\overline{\mathbb{Q}}$ whose $\overline{\mathbb{Q}}$-points parametrizes all pairs $(X,(A,L))$ of a principally polarized abelian variety $(A,L)$ and an irreducible subvariety $X$ of $A$ with $\dim X=r$ and $\deg_{\mathfrak{L}_{g}^{\otimes 4}|_{A}}X=d$.
We have a commutative diagram ((3.7) with $m=1$)
(5.2) ${\mathscr{X}_{\mathbf{H}^{\circ}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{A}_{\mathbf{H}^{\circ}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi}$$\scriptstyle{\iota}$${\mathfrak{A}_{g}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi^{\mathrm{univ}}}$${\mathbf{H}^{\circ}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\iota_{\mathbf{H}^{\circ}}}$${\mathbb{A}_{g}}$
where $\pi|_{\mathscr{X}_{\mathbf{H}^{\circ}}}\colon\mathscr{X}_{\mathbf{H}^{\circ}}\rightarrow\mathbf{H}^{\circ}$ is the universal family. Every object in the diagram (4.2) is a variety defined over $\overline{\mathbb{Q}}$, and every morphism is defined over $\overline{\mathbb{Q}}$. Set ${\mathcal{L}}:=\iota^{*}\mathfrak{L}_{g}^{\otimes 4}$.
@@
There exist constants $c_{2}^{\prime\prime}=c_{2}^{\prime\prime}(r,d,S)>0$ and $c_{3}^{\prime}=c_{3}^{\prime\prime}(r,d,S)>0$ such that the following property holds true. For each $s\in S_{\mathrm{gen}}$, the set
@@
satisfies the following property: $\Sigma_{s}\subseteq X^{\prime}(\overline{\mathbb{Q}})$ for some proper Zariski closed $X^{\prime}\subsetneq\mathscr{X}_{s}$ with $\deg_{{\mathcal{L}}_{s}}(X^{\prime})<c_{2}^{\prime\prime}$.
This proposition is the generalization of the third-named author’s [Küh21, Prop.21] on families of curves, with which the proof of Proposition 5.2 is in line; the main new ingredients are the non-degeneracy construction Proposition 3.4, replacing [Gao20, Thm.1.2’], and Lemma 4.3, which generalizes [DGH21, Lem.6.3] to treat higher-dimensional subvarieties instead of curves.
For readers’ convenience, we divide the proof into 4 steps as in the first-named author’s survey [Gao21, Prop.8.3]. Briefly speaking, the goal of the first 3 steps is to run a family version of the classical approach of Ullmo–Zhang (for the Bogomolov Conjecture) to obtain a generic lower bound on heights closely related to our height of interest $\hat{h}_{{\mathcal{L}}}$. In Step 4, we then deduce Proposition 5.2 for this height bound. Step 4 is similar to the proof of Proposition 4.2, where Lemma 4.3 is applied to get the desired $X^{\prime}$; the main difference of this step and the proof of Proposition 4.2 is that the height inequality of Dimitrov–Gao–Habegger [DGH21, Thm.1.6 and B.1] is replaced by the height bound achieved from the previous steps.
This family version of the Ullmo–Zhang approach differs from the classical one mainly in two aspects. First and naturally, we use a version of equidistribution that applies to families of abelian varieties; for our purpose, the results of the third-named author [Küh21, Thm.1] or the more general results of Yuan–Zhang [YZ21, Thm.6.7] are sufficient replacements of the classical equidistribution result of Szpiro–Ullmo–Zhang [SUZ97]. Second and in contrast to [SUZ97], a new condition, namely non-degeneracy as defined in [DGH21, Defn.B.4], has to be verified for the subvarieties under consideration. For this purpose we apply [Gao20, Thm.10.1]. Notice that in the case of a single abelian variety, every subvariety is non-degenerate so this is a genuinely new aspect of the family case.
@@
Decomposing $S$ into its irreducible components, it suffices to prove the proposition for $S$ irreducible.
From now on, we assume that $S$ is irreducible. We prove the proposition by induction on $\dim S$. The proof of the base step $\dim S=0$ is in fact contained in the induction. We also point out that the induction hypothesis will be applied only in the last step (Step 4).
Step 1 Construct non-degenerate subvarieties.
@@
Let $m_{0}(S^{\prime})>0$ be from Proposition 3.4, and let $m=\max_{S^{\prime}}\{m_{0}(S^{\prime})\}$ with $S^{\prime}$ running over all irreducible components of $\overline{S_{\mathrm{gen}}}$. Then $\mathscr{X}^{[m]}_{S^{\prime}}$ is a non-degenerate subvariety of $\mathcal{A}_{S^{\prime}}^{[m]}$ by Proposition 3.4. Notice that $m=m(S)>0$ depends only on $S$.
By definition of non-degenerate subvariety [Gao21, Defn.6.1], for the Betti form $\omega_{m}$ of $\mathcal{A}_{S^{\prime}}^{[m]}$, there exists a point $\mathbf{x}\in(\mathscr{X}^{[m]}_{S^{\prime}})^{\mathrm{sm}}(\mathbb{C})$ such that
(5.4) $(\omega_{m}|_{\mathscr{X}^{[m]}_{S^{\prime}}(\mathbb{C})}^{\wedge\dim\mathscr{X}^{[m]}_{S^{\prime}}})_{\mathbf{x}}\not=0.$
By generic smoothness, there exists a Zariski open dense subset $U$ of $(S^{\prime})^{\mathrm{sm}}$ such that $\pi^{[m]}|_{\mathscr{X}^{[m],\mathrm{sm}}_{S^{\prime}}\cap(\pi^{[m]})^{-1}(U)}$ is a smooth morphism. We have some freedom to choose $\mathbf{x}$, and we may and do assume $\mathbf{x}$ lies above $U$.
For each $M\geq 1$, recall the proper $S^{\prime}$-morphism
(5.5) $\mathscr{D}_{M}^{\mathcal{A}_{S^{\prime}}^{[m]}}\colon(\mathcal{A}_{S^{\prime}}^{[m]})^{[M+1]}\rightarrow(\mathcal{A}_{S^{\prime}}^{[m]})^{[M]}$
fiberwise defined by $(\mathbf{a}_{0},\mathbf{a}_{1},\ldots,\mathbf{a}_{M})\mapsto(\mathbf{a}_{1}-\mathbf{a}_{0},\ldots,\mathbf{a}_{M}-\mathbf{a}_{0})$, with each $\mathbf{a}_{i}\in\mathcal{A}_{S^{\prime}}^{[m]}(\overline{\mathbb{Q}})$.

### v4

@@ 4.2. The case of principally polarized abelian varieties
Let $\mathbf{H}^{\circ}_{\mathrm{gen}}$ be the subset of $\mathbf{H}^{\circ}(\overline{\mathbb{Q}})$ as defined by (3.9). If the stabilizer $\mathrm{Stab}_{A}(X)$ of $X$ in $A$ has positive dimension, then $X^{\circ}=\emptyset$ by the definition of the Ueno locus and hence the proposition trivially holds true. Therefore, we may and do assume that $\mathrm{Stab}_{A}(X)$ is finite. Then the point $s\in\mathbf{H}^{\circ}(\overline{\mathbb{Q}})$ parameterizing $(X,(A,L))$ is in $\mathbf{H}_{\mathrm{gen}}^{\circ}$.
Invoking Proposition 4.2 with $S=\mathbf{H}^{\circ}$, we obtain constants $c_{1}^{\prime}$, $c_{2}^{\prime}$ and $c_{3}^{\prime}$ depending only on $r$, $8^{r}d$ and $\mathbf{H}^{\circ}$ (and hence ultimately only on $g$, $d$ and $r$) and a proper Zariski closed $X^{\prime}\subsetneq X$ with $\deg_{{\mathcal{L}}|_{A}}(X^{\prime})<c_{2}^{\prime}$ such that the set
$\Sigma_{s}:=\left\{x\in X^{\circ}(\overline{\mathbb{Q}}):\hat{h}_{{\mathcal{L}}|_{A}}(x)\leq c_{1}^{\prime}\max\{1,h_{\overline{\mathbf{H}^{\circ}},{\mathcal{M}}}(s)\}-c_{3}^{\prime}\right\}$
is contained in $X^{\prime}(\overline{\mathbb{Q}})$.
Notice that $\deg_{L}(X^{\prime})<\deg_{{\mathcal{L}}|_{A}}(X^{\prime})$ because of $8c_{1}(L)=c_{1}({\mathcal{L}}|_{A})$ and hence $\deg_{L}(X^{\prime})<c_{2}^{\prime}$.
@@
We are done on replacing $c_{1}^{\prime}$ with $c_{1}^{\prime}/4(c(g)+c^{\prime}(g))$ and $c_{3}^{\prime}$ with $(c_{3}^{\prime}+2c_{1}^{\prime})/4$. ∎
@@
Let $(A,L)$ be a polarized abelian variety with $\deg_{L}A=l$ and let $X$ be an irreducible subvariety generating $A$ and such that $\deg_{L}X=d$.
By Lemma 2.2.(iv), there exists a principally polarized abelian variety $(A_{0},L_{0})$ and an isogeny $u_{0}\colon A\rightarrow A_{0}$ such that $L=u_{0}^{*}L_{0}$; moreover, $\deg(u_{0})=l/g!$. A basic property of the Faltings height [Ray85, Prop. 1.4.1] yields the bound
@@
It is well-known that there exists an isogeny $u\colon A_{0}\rightarrow A$ such that $u_{0}\circ u=[\deg u_{0}]$ on $A_{0}$. So $u^{*}L=(u_{0}\circ u)^{*}L_{0}=[l/g!]^{*}L_{0}$, and $\deg u=(\deg u_{0})^{2g-1}=(l/g!)^{2g-1}$.
Let $X_{0}$ be an irreducible component of $u^{-1}(X)$. Then $u(X_{0})=X$ as $u$ is étale and $X_{0}$ generates $A_{0}$. By the projection formula [Ful98, Prop. 2.5], we have
(4.14) $d^{\prime}:=\deg_{u^{*}L}X_{0}=c_{1}(u^{*}L)^{\dim X}\cdot[X_{0}]\\ =\deg(u|_{X_{0}})\cdot(c_{1}(L)^{\dim X}\cdot[X])\leq\deg(u)\cdot\deg_{L}X=d(l/g!)^{2g-1}.$
It is easy to infer from the definition of the Ueno locus that $X^{\circ}=u(X_{0})^{\circ}\subseteq u(X_{0}^{\circ})$.
Let $c_{0,1}^{\prime}(g,r,d^{\prime})>0$, $c_{0,2}^{\prime}(g,r,d^{\prime})>0$ and $c_{0,3}^{\prime}(g,r,d^{\prime})>0$ be the constants from Proposition 4.4 with $d$ replaced by $d^{\prime}$. Set
$c_{1}^{\prime}:=\min_{1\leq d^{\prime}\leq d(l/g!)^{2g-1}}\{c_{0,1}^{\prime}(g,r,d^{\prime})\}>0,$
$c_{2}^{\prime}:=\max_{1\leq d^{\prime}\leq d(l/g!)^{2g-1}}\{c_{0,2}^{\prime}(g,r,d^{\prime})\}>0,$
and
$c_{3}^{\prime}:=\max_{1\leq d^{\prime}\leq d(l/g!)^{2g-1}}\{c_{0,3}^{\prime}(g,r,d^{\prime})\}>0.$
Then $c_{1}^{\prime}$, $c_{2}^{\prime}$ and $c_{3}^{\prime}$ depend only on $g$, $r$ and $d$. Proposition 4.4 yields the following assertion: the set
$\Sigma_{0}:=\left\{x_{0}\in X_{0}^{\circ}(\overline{\mathbb{Q}}):\hat{h}_{L_{0}\otimes[-1]^{*}L_{0}}(x_{0})\leq c_{1}^{\prime}\max\{1,h_{\mathrm{Fal}}(A_{0})\}-c_{3}^{\prime}\right\}$
is contained in $X_{0}^{\prime}(\overline{\mathbb{Q}})$ for a proper Zariski closed $X_{0}^{\prime}\subsetneq X_{0}$ with $\deg_{L_{0}}(X^{\prime}_{0})\leq c_{2}^{\prime}$.
Let
$\Sigma:=\left\{x\in X^{\circ}(\overline{\mathbb{Q}}):(g!/l)^{2}\hat{h}_{L\otimes[-1]^{*}L}(x)\leq c_{1}^{\prime}\max\{1,h_{\mathrm{Fal}}(A)\}-c_{3}^{\prime}-(1/2)\log(l/g!)\right\}.$
As $X^{\circ}\subseteq u(X_{0}^{\circ})$ and $u^{*}(L\otimes[-1]^{*}L)\cong(L_{0}\otimes[-1]^{*}L_{0})^{\otimes(l/g!)^{2}}$, (4.13) yields $\Sigma\subseteq u(\Sigma_{0})$.
Set $X^{\prime}:=u(X_{0}^{\prime})$. Then $X^{\prime}\subsetneq X$, $\Sigma\subseteq X^{\prime}(\overline{\mathbb{Q}})$, and
$\deg_{L}X^{\prime}=c_{1}(L)^{\dim X^{\prime}}\cdot[X^{\prime}]\leq c_{1}(u^{*}L)^{\dim X^{\prime}}\cdot[X_{0}^{\prime}]=\deg_{u^{*}L}X_{0}^{\prime}\leq c_{2}^{\prime}.$
Replacing $c_{1}^{\prime}$ with $(l/g!)^{2}c_{1}^{\prime}$ and $c_{3}^{\prime}$ with $(l/g!)^{2}(c_{3}^{\prime}+(1/2)\log(l/g!))$, we obtain the assertion of Proposition 4.1. ∎
@@ 5. Applying equidistribution
Let $g$, $l$, $r$ and $d$ be positive integers. There exist constants $c_{2}^{\prime\prime}=c_{2}^{\prime\prime}(g,l,r,d)>0$ and $c_{3}^{\prime\prime}=c_{3}^{\prime\prime}(g,l,r,d)>0$ with the following property. For
  • every polarized abelian variety $(A,L)$ of dimension $g$ defined over $\overline{\mathbb{Q}}$ with $\deg_{L}A=l$, and
  • every irreducible subvariety $X\subseteq A$ generating $A$ with $\dim X=r$ and $\deg_{L}X=d$,
@@
(5.1) $\Sigma:=\left\{x\in X^{\circ}(\overline{\mathbb{Q}}):\hat{h}_{L\otimes[-1]^{*}L}(x)\leq c_{3}^{\prime\prime}\right\}$
is contained in $X^{\prime}(\overline{\mathbb{Q}})$ for a proper Zariski closed $X^{\prime}\subsetneq X$ with $\deg_{L}(X^{\prime})<c_{2}^{\prime\prime}$.
The main idea of the proof this proposition is similar to that of the proof of Proposition 4.1: We put the various pairs of polarized abelian varieties $(A,L)$ and subvarieties $X\subset A$ into finitely many families over the components of the restricted Hilbert scheme constructed in (3.6). The technical core of this section is a related result for each such family (Proposition  5.2). To achieve the uniformity as stated, we need to do however also a careful bookkeeping.
5.1. A preliminary result for the restricted Hilbert scheme
@@
For each $1\leq r\leq g$ and each $d\geq 1$, let $\mathbf{H}^{\circ}:=\mathbf{H}_{r,d}(\mathfrak{A}_{g}/\mathbb{A}_{g})^{\circ}$ be the restricted Hilbert scheme from (3.6); recall that its $\overline{\mathbb{Q}}$-points parametrize all pairs $(X,(A,L))$ of a principally polarized abelian variety $(A,L)$ and an irreducible subvariety $X\subseteq A$ with $\dim X=r$ and $\deg_{\mathfrak{L}_{g}^{\otimes 4}|_{A}}X=d$.
We have a commutative diagram over $\overline{\mathbb{Q}}$ ((3.7) with $m=1$)
[…]
where $\pi|_{\mathscr{X}_{\mathbf{H}^{\circ}}}\colon\mathscr{X}_{\mathbf{H}^{\circ}}\rightarrow\mathbf{H}^{\circ}$ is the universal family. Set ${\mathcal{L}}:=\iota^{*}\mathfrak{L}_{g}^{\otimes 4}$.
@@
Let $S\subseteq\mathbf{H}^{\circ}$ be a (not necessarily irreducible) subvariety and let $S_{\mathrm{gen}}\subseteq S(\overline{\mathbb{Q}})$ be the subset defined by (3.9).
There exist constants $c_{2}^{\prime\prime}=c_{2}^{\prime\prime}(r,d,S)>0$ and $c_{3}^{\prime}=c_{3}^{\prime\prime}(r,d,S)>0$ such that the following property holds true. For each $s\in S_{\mathrm{gen}}(\overline{\mathbb{Q}})$, the set
@@
is contained in a proper Zariski closed subset $X^{\prime}\subsetneq\mathscr{X}_{s}$ with $\deg_{{\mathcal{L}}_{s}}(X^{\prime})<c_{2}^{\prime\prime}$.
This proposition is a generalization of [Küh21, Prop. 21] on families of curves. For convenience of the reader, we divide the proof into five steps. The goal of the first four steps is to run a family version of the classical approach of Ullmo [Ull98] and Zhang [Zha98] to the Bogomolov conjecture in order to obtain a generic lower bound on heights that are closely related to $\hat{h}_{{\mathcal{L}}}$, which is our height of interest. In the final step, we then deduce Proposition 5.2 from this height bound. As in the proof of Proposition 4.2, we use Lemma 4.3 to deal with the non-generic points for which we do not obtain a height bound; this gives rise to the exceptional set $X^{\prime}$.
This family version of the Ullmo–Zhang approach differs from the classical one mainly in two aspects. First, a version of equidistribution is used that applies to families of abelian varieties; for our purpose, [Küh21, Thm. 1] or the more general results [YZ26, Thm. 6.7] and [Gau26] are sufficient replacements of the classical equidistribution result of Szpiro–Ullmo–Zhang [SUZ97]. Second and in contrast to [SUZ97], a new condition, namely non-degeneracy as defined in [DGH21, Defn. B.4], has to be verified for the subvarieties under consideration. Notice that every subvariety is non-degenerate in the case of a single abelian variety. So this extra verification is a genuinely new aspect of working in families and in practice this step is often very difficult.
@@
Decomposing $S$ into its irreducible components, it suffices to prove the proposition for irreducible $S$.
We prove the proposition by induction on $\dim S$. The base case $\dim S=0$ is in fact contained in the first three steps of the following proof. In contrast to Proposition 4.2 it is a highly non-trivial statement, namely the classical Bogomolov conjecture. One could alternatively cite [Ull98, Zha98] for the base case and hence we do not state its proof separately as an induction start here. The induction hypothesis is only utilized in Step 5 and is unnecessary in case $\dim S=0$.
Step 1: Construction of non-degenerate subvarieties
@@
Let $m_{0}(S^{\prime})>0$ be from Proposition 3.4, and let $m=\max_{S^{\prime}}\{m_{0}(S^{\prime})\}$ with $S^{\prime}$ running over all irreducible components of $\overline{S_{\mathrm{gen}}}$. Then $\mathscr{X}^{[m]}_{S^{\prime}}$ is a non-degenerate subvariety of $\mathcal{A}_{S^{\prime}}^{[m]}$ by Proposition 3.4. Notice that $m>0$ depends only on $S$. Let us write $\omega_{m}$ for the Betti form on $\mathcal{A}_{S^{\prime}}^{[m]}$. By [DGH21, Prop. 2.2.(iii)], there exists a point $\mathbf{z}\in(\mathscr{X}^{[m]}_{S^{\prime}})^{\mathrm{sm}}(\mathbb{C})$ such that
(5.4) $(\omega_{m}|_{\mathscr{X}^{[m]}_{S^{\prime}}(\mathbb{C})}^{\wedge(rm+\dim(S^{\prime}))})_{\mathbf{z}}\not=0.$
Since this condition on $\mathbf{z}$ is open (in the usual topology), we may and do assume that $\mathbf{z}$ lies over a smooth point of $S^{\prime}$.
For each $M\geq 1$, we consider the proper $S^{\prime}$-morphism
(5.5) $\mathscr{D}_{0}\colon\ \$$(\mathcal{A}_{S^{\prime}}^{[m]})^{[M+1]}$ $\longrightarrow$ $(\mathcal{A}_{S^{\prime}}^{[m]})^{[M]}$
$(\mathbf{a}_{0},\mathbf{a}_{1},\ldots,\mathbf{a}_{M})$ $\longmapsto$ $(\mathbf{a}_{1}-\mathbf{a}_{0},\ldots,\mathbf{a}_{M}-\mathbf{a}_{0}),$
using the fiberwise group structure.